OpenStudy (anonymous):

I need help finding two irrational solutions for g(x)=x^2+6x+1

OpenStudy (campbell_st):

well use the general quadratic formula \[x = \frac{-b \pm \sqrt{b^2 - 4ac}}{2a}\] in your question you have a = 1, b = 6 and c = 1 just substitute and simplify for the irrational roots

OpenStudy (anonymous):

so, \[x=\frac{ -6\pm \sqrt{6^{2-4(1)(1)}} }{ 2(1) }\]

OpenStudy (campbell_st):

well nearly \[x = \frac{-6 \pm \sqrt{6^2 - 4 \times 1 \times 1}}{2 \times 1}\]